### v1

# GFlowGR: Fine-tuning Generative Recommendation Frameworks with Generative Flow Networks

Abstract.
Generative recommendations (GR), which usually include item tokenizers and generative Large Language Models (LLMs), have demonstrated remarkable success across a wide range of scenarios. The majority of existing research efforts primarily concentrate on developing powerful item tokenizers or advancing LLM decoding strategies to attain superior performance. However, the critical fine-tuning step in GR frameworks, which is essential for adapting LLMs to recommendation data, remains largely unexplored. Current approaches predominantly rely on either the next-token prediction loss of supervised fine-tuning (SFT) or recommendation-specific direct preference optimization (DPO) strategies. Both methods ignore the exploration of possible positive unobserved samples, which is commonly referred to as the exposure bias problem. To mitigate this problem, this paper treats the GR as a multi-step generation task and constructs a GFlowNets-based fine-tuning framework (GFlowGR). The proposed framework integrates collaborative knowledge from traditional recommender systems to create an adaptive trajectory sampler and a comprehensive reward model. Leveraging the diverse generation property of GFlowNets, along with sampling and heuristic weighting techniques, GFlowGR emerges as a promising approach to mitigate the exposure bias problem. Extensive empirical results on two real-world datasets and with two different GR backbones highlight the effectiveness and robustness of GFlowGR.
Recommender Systems; Generative Recommendations
$\dagger$ Work was conducted during the internship of Yejing Wang at Alibaba.
* Xiangyu Zhao is corresponding author.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Large Language Models (LLMs) possess exceptional capabilities in understanding, reasoning, and generation. As a result, they are increasingly utilized across various domains, including healthcare, education, and science (Naveed et al., 2023), as well as for information retrieval and recommendations in numerous online services, such as E-commerce (Li et al., 2024a; Zhai, 2024). Notably, the impressive generative capabilities of LLMs have spurred the development of generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023). As visualized in Figure 1 (a), these frameworks typically consist of three key components (Li et al., 2024b): item tokenizers, user prompts, and generative LLMs. Meanwhile, the training paradigms usually consist of two stages: optimizing the item tokenizer to represent items with several tokens and fine-tuning the LLMs for recommendation objectives like next-item prediction.
Existing research on GR has primarily concentrated on enhancing the item tokenizer (Zheng et al., 2025; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024) and the associated optimization processes, designing user prompts (Liu et al., 2024a; Wang et al., 2024b), and advancing generative LLMs (Lin et al., 2025c; Ding et al., 2024). However, there has been little focus on innovating the fine-tuning step of LLMs. While this step is crucial for GR frameworks: the semantics of recommendations can significantly differ from those of LLMs pretrained for general language purposes (Lin et al., 2025a; Wu et al., 2024). For instance, terms like “Comment” represent specific user behaviors within recommendation data, even though they are commonly used verbs in general language tasks. Moreover, the item identifier in GR may consist of out-of-vocabulary tokens for pretrained LLMs (Huang et al., 2024; Wang et al., 2024a; Rajput et al., 2023), which are unlikely to be generated by zero-shot LLMs.
Current LLM fine-tuning techniques applied for GR can be generally divided into two categories (Wu et al., 2025; Wang et al., [n. d.]): supervised fine-tuning (SFT) and reinforcement learning-based fine-tuning (RLFT). The former optimizes LLMs with the next token prediction loss (Brown et al., 2020; Radford et al., 2019; Hu et al., 2022) and handcrafted auxiliary alignment loss (Liu et al., 2024a; Zheng et al., 2024; Wang et al., 2024b), which focuses solely on positive samples within the dataset. On the one hand, SFT does not incorporate any negative samples for comparison with the positive ones, leaving a gap with the ranking objective of recommender systems (Chen et al., [n. d.]). On the other hand, SFT cannot explore potential positive samples to extend the collected data, which is always sparse (Gao et al., 2022). With the exploration ability, RLFT methods like proximal policy optimization (PPO) (Schulman et al., 2017), and direct preference optimization (DPO) (Rafailov et al., 2023) can improve the model capacity on the basis of SFT by achieving the maximum reward. For example, recent DPO-based methods facilitate LLM fine-tuning by integrating augmented negative samples for GR (Chen et al., [n. d.]; Bai et al., 2024; Liao et al., 2024; Gao et al., 2025a; Deng et al., 2025), enabling LLMs to rank candidate items. However, both solutions cannot match the diverse nature of recommender systems that users may prefer multiple items, while only one among them can be observed. SFT regards the unobserved interactions as negative, and RLFT methods predefine the largest reward to the observed samples. This reliance on the collected data that fails to fully reveal the real user preference can lead to biased user preferences, known as exposure bias (Chen et al., 2023). Therefore, it calls for an LLM fine-tuning solution that can explore potential positive samples and leverage augmented samples to achieve diverse results.
Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: by training LLMs to generate token sequences with probabilities proportional to a specific reward distribution (Bengio et al., 2023), GFlowNets can identify potential positive candidates and thus extend the recommendation diversity. This capability is supported by recent studies demonstrating the application of GFlowNets for multi-step reasoning tasks (Kang et al., 2025; Takase et al., 2024; Yu et al., 2024) and compositional generation (Jain et al., 2022). Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token.However, applying GFlowNets for GR still faces significant challenges: how to organize training trajectories (samples) and establish a reward model to assess the terminal state for GFlowNets. Specifically, the collected data consists solely of positive interaction logs from users, resulting in sparse training trajectories for GFlowNets. This scenario necessitates high-quality augmented trajectories for robust training. Furthermore, merely treating these augmented trajectories as negative samples and assigning them low rewards is inadequate for overcoming the exposure bias problem, as these augmented trajectories are not consistently negative.
To address these limitations, this paper constructs the GFlowNets-based fine-tuning paradigm for GR (GFlowGR). Drawing inspiration from prior insights on aligning LLMs with collaborative recommendation models (Hong et al., [n. d.]; Yang et al., 2024; Lin et al., 2025a), we incorporate collaborative scores when designing the trajectory sampling and reward model for GFlowGR. Specifically, we develop an adaptive trajectory sampler that generates easy-to-hard augmented trajectories, thereby implementing a curriculum-style training approach (Wang et al., 2021a) to consistently provide high-quality training data. Additionally, we construct a multi-faceted reward system that thoroughly considers augmentation signals, collaborative scores, and semantic similarities to effectively differentiate possibly preferred items from the sampled unobserved trajectories. Finally, to further mitigate the exposure bias problem, we assign confidence weights to each sample based on the agreement between augmentation signals and collaborative scores. We summarize the major contributions of this paper as follows:
  • We formulate GR as a multi-step generation process and develop a GFlowNets-based fine-tuning framework, GFlowGR. This framework helps mitigate the exposure bias problem by learning generation probabilities that are proportional to the reward.
  • We implement an adaptive trajectory sampler to select high-quality training samples for GFlowGR and establish a comprehensive reward model to consider various reward signals effectively.
  • Our extensive experimental results, obtained from two public benchmarks using two backbone GR models, demonstrate the effectiveness and versatility of GFlowGR.
Figure 1. Overview of GR and GFlowNets, with the length of item identifiers $L=3$.
2. Preliminary
Generative Recommendation (GR). In this paper, we define GR frameworks by referencing recent works (Rajput et al., 2023; Wang et al., 2024a). A typical GR workflow is illustrated in Figure 1 (a). The framework first constructs an item tokenizer using item features to decompose items $v$ into token sequences $\{t_{l}\}_{l<L}$, serving as item identifiers. The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence $<a\_1><b\_4><c\_2>$ based on its title and descriptions. To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $<\underline{X}>$). These prompts, exemplified by tokenized interaction histories (blue tokens for jackets, shirts, pants) and profiles, are fed into LLMs to generate token sequences representing the most probable next item. As shown in the upper portion of Figure 1 (a), this process produces $<a\_2><b\_4><c\_10>$, decoding to a blue polo shirt as the recommended item.
To achieve the recommendation goal, the training paradigm of GR comprises two stages: optimizing the item tokenizer ($\mathcal{L}_{T}$) and the generative LLMs ($\mathcal{L}_{\text{GR}}$), respectively. For the tokenizer, one prevalent tokenization strategy is based on RQ-VAE solutions (Lee et al., 2022; Rajput et al., 2023), where $\mathcal{L}_{T}$ includes a codebook loss and a reconstruction loss. Studies enhancing the tokenizer often incorporate auxiliary losses into $\mathcal{L}_{T}$, such as diversity (Wang et al., 2024a) and collaborative-constrained losses (Lin et al., 2025b). $\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::::
(1) $\mathcal{L}_{\text{GR}}(U,v)=-\sum_{l=1}^{L}\mathrm{log}\mathbb{P}_{\text{GR}}(t_{l}|U,\{t_{i}\}_{i<l})$
where $L$ denotes the length of item identifiers (e.g., $L=3$ for the visualized examples in this paper). We denote the probability that LLMs generate the next ground-truth token $t_{l}$ given the prefix $U,\{t_{i}\}_{i<l}$ as $\mathbb{P}\left(t_{l}\,\big{|}\,U,\{t_{i}\}_{i<l}\right)$. Here, $\{t_{i}\}_{i<l}$ represents the item identifier tokens preceding $t_{l}$.
Generative Flow Networks (GFLowNets). The core idea of Generative Flow Networks (GFlowNets) stems from the generation of compositional objects through a sequence of stochastic steps. We formally define the state transition sequence as $\tau=(s_{0}\to s_{1}\to\dots\to s_{L})$, where $s_{L}$ represents the terminal generation state and $\tau$ denotes a specific generation trajectory. The training objective of GFlowNets is to align the generation probability of any trajectory $\tau$ with the reward obtained at its terminal state, such that $\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$  (Bengio et al., 2021; Takase et al., 2024). To achieve this, GFlowNets introduce a flow estimator $\mathcal{F}(s_{l})$ to model the likelihood of traversing state $s_{l}$, transforming the proportionality objective into a flow-matching principle: the incoming flow to a state must equal the outgoing flow. This is formalized as:
(2) $\mathcal{F}(s_{l})\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathcal{F}(s_{l+1})\mathbb{P}_{B}(s_{l}|s_{l+1})$
Here, $\mathbb{P}_{F}(s_{l+1}|s_{l})$ denotes the forward transition probability from state $s_{l}$ to $s_{l+1}$, while $\mathbb{P}_{B}(s_{l}|s_{l+1})$ represents the backward probability that state $s_{l+1}$ originated from $s_{l}$. This bidirectional flow balance ensures that trajectories with higher terminal rewards $\mathcal{R}(s_{L})$ are assigned greater generation probabilities, enabling efficient exploration of complex combinatorial spaces (Bengio et al., 2023).
To facilitate practical implementation, two optimization-friendly variants are derived from the flow-matching principle (Malkin et al., 2022; Bengio et al., 2021): detailed balance (DB) and trajectory balance (TB). The former focuses on stepwise flow equilibrium, while the latter enforces balance at the full-trajectory level. Their respective loss functions are defined:
(3) $\mathcal{L}_{\text{DB}}(\tau)=\sum_{l=0}^{L-1}\Big{(}\mathrm{log}\frac{\mathcal{F}(s_{l})\mathbb{P}_{F}(s_{l+1}|s_{l})}{\mathcal{F}(s_{l+1})\mathbb{P}_{B}(s_{l}|s_{l+1})}\Big{)}^{2}$
(4) $\mathcal{L}_{\text{TB}}(\tau)=\Big{(}\mathrm{log}\frac{\mathcal{Z}\prod_{l=0}^{L-1}\mathbb{P}_{F}(s_{l+1}|s_{l})}{\mathcal{R}(s_{L})\prod_{l=0}^{L-1}\mathbb{P}_{B}(s_{l}|s_{l+1})}\Big{)}^{2}$
Here, $\mathcal{Z}$ is a global scalar that approximates $\mathcal{F}(s_{0})$ (Malkin et al., 2022).
Motivation. Two key insights motivate us to apply GFlowNets for GR: first, GR inherently involves a multi-step generation process well-suited for GFlowNet modeling. As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., $<a\_2>$, $<b\_4>$, $<c\_10>$) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo). This sequential structure aligns naturally with GFlowNets’ trajectory-based framework. Second, GFlowNets offer two critical advantages for GR: (a) Diverse High-Reward Generation: by linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high predicted rewards. Theoretical analyses (Bengio et al., 2021, 2023) and applications in domains like DNA design (Jain et al., 2022) and multi-step reasoning (Kang et al., 2025; Yu et al., 2024) validate this capability. In recommendation, this addresses exposure bias by treating unobserved interactions as potential positives rather than default negatives. Particularly, during training, GFlowNets can identify augmented items with high reward scores as potential positive and prioritize them over low-reward alternatives. (b) Off-Policy Data Utilization: GFlowNets enable off-policy learning, allowing GR frameworks to incorporate augmented interactions into training. This extends SFT, which relies solely on collected positive samples, to leverage a broader range of potential user preferences. The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards, GFlowNets empower GR to prioritize the unseen blue jacket (aligned with the user’s formal dressing history) over the yellow undershirt, even when both are unobserved. This contrasts with conventional methods that treat unobserved items equally, thereby enhancing diversity and mitigating exposure bias.
Implementing GFlowNets for GR involves straightforward computation of forward and backward probabilities: the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})$ corresponds to the LLM’s token generation probability $\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$; the backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ due to the sequential nature of token generation ($s_{l+1}=\{t_{i}\}_{i\leq(l+1)}=\{s_{l},t_{l+1}\}$, $s_{l+1}$ is uniquely derived from $s_{l}$). Thus, the core design challenges reduce to constructing training trajectories $\tau$ and designing the reward model $\mathcal{R}(s_{L})$. These components are critical for GFlowNets and will be addressed in subsequent sections.
3. Methodology
In this section, we detail the proposed GFlowGR framework, including its adaptive trajectory sampler, comprehensive reward model, and training/inference procedures. The advantages of GFlowGR in addressing exposure bias and negative sampling are summarized in the Supplementary Material.
The architecture of GFlowGR is depicted within the dashed block of Figure 2. Notably, GFlowGR modifies only the LLM fine-tuning stage for GR while keeping the inference pipeline unchanged. The framework integrates a collaborative model (CM) that is capable of scoring candidate items (blue dotted block) based on user interaction histories (pink dotted block) as a core component of both the trajectory sampler and reward model. The overall workflow is as follows: for a given input pair $(U,v)$, the adaptive trajectory sampler selects augmented samples $v^{\prime}$ as extended generation trajectories. GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for both the original trajectory $\tau$ (corresponding to the ground-truth $v$) and augmented trajectories $\tau^{\prime}$ (corresponding to $v^{\prime}$). The reward model $\mathcal{R}$ integrates collaborative scores ($r_{c}$), augmentation signals from the trajectory sampler ($r_{a}$), and token similarities between positive and augmented item identifiers ($r_{\text{sim}}$) to compute rewards. Finally, GFlowGR computes the GFlowNet loss $\mathcal{L}_{\text{GFN}}$ for the LLM fine-tuning.
Figure 2. Framework of GFlowGR. The visualized example shows a positive sample (blue polo shirt) and an augmented sample (yellow undershirt). Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity.
3.1. Adaptive Trajectory Sampler
The sparsity of recommendation data is a fundamental challenge, as only a minimal fraction of user-item interactions are observed (Gao et al., 2022). Consequently, constructing training trajectories solely from collected data is insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset. A naive approach to address this is through data augmentation (Dang et al., 2024) or random negative sampling (Ma et al., 2024)—for instance, randomly selecting $K$ augmented samples $\{v^{\prime}_{k}\}_{k\leq K}$ such that $v\neq v^{\prime}_{k}$ for all $k$ (Yang et al., 2020; Wang et al., 2021b). However, such randomness introduces training instability and often yields samples misaligned with the model’s evolving learning dynamics. For example, an obviously irrelevant augmented sample that was effective as a negative in early training epochs may become uninformative in later stages, as the model learns to easily distinguish it from true positives.
To address this, we propose an adaptive trajectory sampler that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner. We assume that during the early training phase, LLMs struggle to distinguish between positive and augmented samples, while in later stages, overly simple samples contribute minimally to model improvement (Wang et al., 2021a). Inspired by recent work demonstrating the benefits of integrating collaborative knowledge into LLM-based recommendation systems (Lin et al., 2025a), we leverage a pretrained collaborative model (CM) to guide sample augmentation. Specifically, we score $N$ candidate augmentations $\{v^{\prime}_{n}\}_{n\leq N}$ using $\hat{y}_{CM}^{v^{\prime}_{n}}=\mathrm{CM}(v^{\prime}_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v^{\prime}_{n}}$ represents the collaborative score of item $v^{\prime}_{n}$ to user $U$.
During initial training epochs, we prioritize sampling items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire. As training progresses, the sampler gradually incorporates harder augmented samples with higher collaborative scores—for example, the blue jacket in the visualized example, which aligns more closely with the user’s preferences but remains unobserved in historical data. This curriculum-based approach enables the sampler to adaptively select trajectories rooted in collaborative knowledge, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
3.2. Comprehensive Reward Model
The reward model is pivotal in GFlowNets, as it directly aligns generation probabilities with task-specific objectives (Bengio et al., 2021, 2023). In recommendation contexts, prior GFlowNet reward models typically rely on observed user interactions (e.g., item ratings from 1–6 (Liu et al., 2024b, 2023)), but this approach fails for unobserved samples. Assigning zero rewards to augmented samples may ignore the latent relevance and exacerbate exposure bias (Chen et al., 2023).
To address this, we introduce a comprehensive reward model that integrates three complementary signals: augmentation status, collaborative scores, and token similarity. For simplicity, we describe the framework using a single augmented sample ($K=1$, easily extendable to multiple samples). Let $\tau=\{s_{l}\}_{l\leq L}$ denote the trajectory of the observed positive item $v$, and $\tau^{\prime}=\{s^{\prime}_{l}\}_{l\leq L}$ denote the trajectory of an augmented item $v^{\prime}$. Three aspects of reward are defined as:
  • Augmentation Signal ($r_{a}$): marks whether a sample is observed (positive) or augmented, i.e., $r_{a}=1$ for the ground-truth trajectory $\tau$ and $r^{\prime}_{a}=0$ for the augmented trajectory $\tau^{\prime}$.
  • Collaborative Score ($r_{c}$): the collaborative score from the same model, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$.
  • Token Similarity ($r_{\text{sim}}$): rewards partial correctness by measuring the number of shared tokens with the positive sample. Specifically, for $\tau^{\prime}$, $r^{\prime}_{\text{sim}}=\sum_{l=1}^{L}\mathbb{I}(t_{l}=t^{\prime}_{l})$, where $\mathbb{I}(\cdot)$ is an indicator function. And $r_{\text{sim}}=L$ for $\tau$. For the visualized example, $r^{\prime}_{\text{sim}}=1$, as it has one shared token with the ground truth.
We sum up these reward signals as the final reward for each trajectory, which can be formally expressed as $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$.
3.3. Training and Inference
Training. The training objective of GFlowGR integrates two complementary loss components: the vanilla next-token prediction loss $\mathcal{L}_{\text{GR}}$ for positive samples, the GFlowNet training loss $\mathcal{L}_{\text{GFN}}$ over sampled trajectories. The inclusion of $\mathcal{L}_{\text{GR}}$ (Equation (1)) ensures the model retains foundational recommendation capabilities while adapting the LLM to GR tasks. For $\mathcal{L}_{\text{GFN}}$, substituting the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ and backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ into Equation (3) and (4) yields:
(5) $\mathcal{L}_{\text{DB}}(\tau)=\sum_{l=0}^{L-1}\Big{(}\mathrm{log}\frac{\mathcal{F}(s_{l})\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})}{\mathcal{F}(s_{l+1})}\Big{)}^{2}$
(6) $\mathcal{L}_{\text{TB}}(\tau)=\Big{(}\mathrm{log}\frac{\mathcal{Z}\prod_{l=0}^{L-1}\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})}{\mathcal{R}(s_{L})}\Big{)}^{2}$
$\mathcal{F}$, and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5). $\mathcal{L}_{\text{GFN}}$ can be instantiated using either $\mathcal{L}_{\text{DB}}$ or $\mathcal{L}_{\text{TB}}$. The complete loss function for $(U,v,\tau)$ is:
(7) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\mathcal{L}_{\text{GFN}}(\kappa)$
where $\kappa$ represents either the positive trajectory $\tau$ or an augmented trajectory $\tau^{\prime}$, and $\lambda$ is the hyperparameter to balance different objectives. To mitigate exposure bias, we weight each trajectory by the inverse disagreement between the augmentation signal $r_{a}$ and the collaborative score $r_{c}$:
(8) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\frac{1}{(r_{a}-r_{c})^{2}}\mathcal{L}_{\text{GFN}}(\kappa)$
This heuristic weighting strategy downweights examples where the collaborative model exhibits high uncertainty (large $|r_{a}-r_{c}|$), which often correspond to hard negatives or ambiguous interactions (Chen et al., 2023).
Table 1. Overall performance of baselines and GFlowGR. The boldface refers to beating all baselines. “*” indicates the statistically significant improvements (i.e., one-sided t-test with $p<0.05$) over the best baseline. For all metrics, the higher is better.
Model Finetuning Beauty Yelp
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0204 0.0334 0.0132 0.0174
PPO 0.0020 0.0032 0.0013 0.0017 0.0004 0.0005 0.0003 0.0003
GRPO 0.0347 0.0543 0.0229 0.0291 0.0193 0.0328 0.0126 0.0169
DPO 0.0308 0.0483 0.0200 0.0256 0.0100 0.0182 0.0060 0.0087
S-DPO 0.0314 0.0492 0.0208 0.0265 0.0212 0.0333 0.0137 0.0176
SPRec 0.0312 0.0487 0.0205 0.0258 0.0103 0.0186 0.0067 0.0090
IPA 0.0306 0.0481 0.0198 0.0245 0.0101 0.0179 0.0066 0.0093
GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0221 0.0345 0.0145 0.0185
GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0256* 0.0404* 0.0166* 0.0213*
LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0205 0.0341 0.0131 0.0175
PPO 0.0128 0.0206 0.0087 0.0112 0.0002 0.0004 0.0001 0.0002
GRPO 0.0384 0.0592 0.0251 0.0317 0.0228 0.0380 0.0146 0.0195
DPO 0.0358 0.0571 0.0238 0.0307 0.0172 0.0286 0.0115 0.0151
S-DPO 0.0390 0.0607 0.0259 0.0336 0.0247 0.0389 0.0161 0.0206
SPRec 0.0370 0.0573 0.0249 0.0315 0.0190 0.0303 0.0124 0.0160
IPA 0.0367 0.0586 0.0252 0.0320 0.0197 0.0293 0.0129 0.0163
GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0252 0.0399 0.0166 0.0213
GFlowGR-TB 0.0417* 0.0632* 0.0276* 0.0345* 0.0261* 0.0413* 0.0173* 0.0222*
Inference. The inference process remains unchanged from standard GR frameworks: user prompts are constructed as before, and the fine-tuned LLM generates item identifiers sequentially. By design, GFlowGR enhances GR without altering the inference pipeline, ensuring seamless integration into existing systems.
4. Experiments
4.1. Experimental Setups
Datasets. We used two real-world datasets to evaluate GFlowGR, namely Yelp and Amazon Beauty. The data processing and split are following previous GR works (Wang et al., 2024a; Zheng et al., 2024).
Baselines and Backbones. We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: PPO (Schulman et al., 2017), GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n. d.]), SPRec (Gao et al., 2025a), IPA (Deng et al., 2025). Notably, the specially designed auxiliary loss in addition to SFT is regarded as a part of the GR backbones. And we install GFlowGR to two backbones: TIGER (Rajput et al., 2023) and LETTER (Wang et al., 2024a).
Evaluation Metrics. We evaluate all methods using two metrics, namely Hit Rate (R) and Normalized Discounted Cumulative Gain (NDCG), on top-5 and top-10 recommended lists, yielding four metrics per dataset: R@5, R@10, N@5, and N@10. To ensure robust results, each test is repeated three times with different random seeds ({42, 43, 44}).
4.2. Main Results
To evaluate the effectiveness of GFlowGR, we tested its two variants (DB and TB) against various fine-tuning methods. The results are summarized in Table 1, which consistently demonstrates that GFlowGR outperforms baseline approaches. Below are detailed observations:
Comparison with SFT. Both GFlowGR variants consistently outperform SFT across all datasets and metrics, demonstrating that GFlowGR effectively enhances model capacity through augmented samples, flow-matching objectives, and integrated collaborative knowledge. Notably, GFlowGR-TB achieves more significant improvements than GFlowGR-DB, particularly on Yelp with TIGER, likely because GR’s intermediate states do not uniquely identify items, making whole-trajectory optimization (via TB’s trajectory-level balance) more effective by leveraging one-to-one trajectory-item correspondences.
Comparison with On-policy RLFT. On-policy strategies (especially reward-based PPO) fail to achieve competitive results for GR. This may stem from out-of-vocabulary item identifiers: LLMs pretrained on general language tasks struggle to explore the OOV item space without a well-defined reward model, leading to poor PPO performance. GRPO alleviates this dependency on the reward model via relative advantages. (Shao et al., 2024)
Comparison with Off-policy RLFT. Vanilla DPO (Rafailov et al., 2023) fails to stably improve upon SFT, particularly on large-scale Yelp data, due to its use of fixed and randomly sampled “Rejected” items. In contrast, SPRec (Gao et al., 2025a) (using self-play sampling) and IPA (Deng et al., 2025) (leveraging collaborative knowledge) enhance DPO through dynamic, guided negative sampling. Among DPO-based methods, S-DPO (Chen et al., [n. d.]), which is capable of incorporating multiple negative samples, achieves the best results and competes closely with GFlowGR-DB.
Figure 3. Parameter study on $K$ and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to $K$.
4.3. Parameter Analysis
In this section, we analyze GFlowGR ’s parameter sensitivity, focusing on the number of augmented trajectories $K\in\{0,1,3,5\}$ and the balancing weight $\lambda\in\{0.01,0.1,1,10,100\}$ in Equation (8). Figure 3 visualizes results for TIGER on the Beauty dataset. As $K$ and $\lambda$ increase, performance follows a reverse ‘V’ curve: For $K$, small values (e.g., $K=0$) fail to provide sufficient diversity for sparse data, while large $K=5$ introduces noisy augmented samples that mislead training. For $\lambda$, a moderate weight of $1.0$ optimizes $\mathcal{L}_{\text{GFN}}$, balancing flow-matching objectives with the core recommendation task. Overemphasizing $\mathcal{L}_{\text{GFN}}$ (e.g., $\lambda\geq 10$) degrades GR performance, highlighting the critical role of SFT in maintaining basic recommendation capabilities.
4.4. Ablation Status
In this section, we validate the critical components of GFlowGR: the curriculum-style adaptive trajectory sampler, comprehensive reward model, and uncertainty-based trajectory reweighting. Table 2 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of adaptive sampling (w/o Ada), omission of Augmented Trajectories (w/o Traj), and removal of reweighting (w/o Wt). Key findings include: (a) Loss of any reward signal (augmentation signals, collaborative scores, or token similarities) degrades performance, confirming the necessity of the comprehensive reward model. Notably, $r_{a}$ has the most significant impact, while $r_{c}$ contributes marginally due to CM producing less accurate predictions than GR (Rajput et al., 2023; Wang et al., 2024a), leaving to noisy $r_{c}$ signals. (b) Removing adaptive sampling or using random sampling reduces effectiveness due to unstable training or low-quality augmented samples. (c) Uncertainty-based reweighting contributes to alleviating exposure bias and improving model robustness, validating its role in stabilizing training.
Table 2. Ablation study with TIGER on Beauty.
Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj w/o Wt GFlowGR
R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0388 0.0407
R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0615 0.0651
N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0254 0.0266
N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0327 0.0344
4.5. In-depth Analysis
Case Study. We visualize a case in Figure 4 to study GFlowGR’s potential in addressing exposure bias. The example shows GFlowGR recommending a diverse set of items, including ‘Shampoo’, ‘Scrub’, and ‘Cleanser’ with the ‘Tomatoes’ flavor (bold green), which is a new taste for the user. While SFT focuses on ‘Carrots’-flavored ‘Shampoo’ and ‘Conditioner’ from the same brand, which the user recently purchased (bold orange and red). This suggests GFlowGR alleviates exposure bias by exploring diverse trajectories, generating exploratory results with different flavors and product.
Diversity. For this experiment, we analyze diversity in top-10 recommendations from GFlowGR and SFT across two dimensions. (a) Generation Distribution: We compare models’ item recommendation probabilities against a normal distribution using KL divergence (rather than a uniform distribution because item popularities are inherently different for recommendations (Chen et al., 2023)). GFlowGR achieves a closer match (KL=3.13 vs. SFT’s 3.24). In addition, the generation probability over all items has a 2.2% smaller standard deviation with GFlowGR, indicating a smoother, more diverse distribution. (b) Collaborative Scores: We calculate the collaborative score for the recommended items. The result shows that GFlowGR’s recommended items have 0.4% lower average collaborative scores but a broader range (0.45 vs. SFT’s 0.42) and 3.6% higher standard deviation, showing it prioritizes non-mainstream items while offering greater score diversity. These results demonstrate GFlowGR’s enhanced ability to balance relevance and novelty compared to SFT.
Figure 4. Case study on user-2089 from Beauty with TIGER: We visualized three items from the top-10 list of each method for the brand ‘Yes to’, where the top-1 result of GFlowGR is correct.
5. Related Work
Generative Recommendations. Existing GR research can be categorized into three directions: pipeline design, item tokenizer innovations, and LLM decoding enhancement. For pipeline design, TIGER (Rajput et al., 2023) pioneered out-of-vocabulary item tokenization and GR workflows, while LC-Rec (Zheng et al., 2024) and ColaRec (Wang et al., 2024b) enhance alignment via tasks or graph collaborative models, with extensions to multi-behavior scenarios (Liu et al., 2024a). The second category focuses on plug-in tokenizers (e.g., LETTER (Wang et al., 2024a) for collaborative embeddings) to boost arbitrary GR frameworks. The third category optimizes decoding (Lin et al., 2025c; Ding et al., 2024). However, LLM fine-tuning for GR remains underexplored, with DPO-based IPA (Deng et al., 2025) as a rare example. The proposed GFlowGR is the first to study the exposure bias and diversity.
Generative Flow Network Applications. GFlowNets are designed for compositional generation to produce diverse, high-reward outputs (Bengio et al., 2021, 2023), with applications spanning biology (Jain et al., 2022), scientific discovery (Jain et al., 2023), and multi-step reasoning (Yu et al., 2024; Kang et al., 2025). While prior work has applied GFlowNets to discriminative recommendation tasks like sequential recommendation (Liu et al., 2024b) and list-wise recommendation (Liu et al., 2023), as well as language-based recommendations (Gao et al., 2025b), we are the first to explore GFlowNets for generative recommendations.
6. Conclusion
In this paper, we propose GFlowGR, a GFlowNets-based fine-tuning framework for GR. Specifically, we design an adaptive trajectory sampler that leverages collaborative knowledge to select augmented trajectories, enabling the model to explore diverse item spaces. We also introduce a comprehensive reward model that integrates multi-faceted signals, including augmentation status, collaborative scores, and token similarities, to provide nuanced feedback for trajectory evaluation. During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives, with trajectory-specific losses reweighted based on uncertainty. By adaptive sampling, heuristic reweighting, and flow-matching targets, GFlowGR enhances recommendation diversity and addresses the exposure bias challenge inherent in GR frameworks. Extensive experimental results validate the effectiveness of our approach across multiple datasets, backbones, and metrics.